A company received an invoice for $3450 dated March 23, terms 2/10.

What payment must be made on March 31 to reduce the debt by $1500?




1
Expert's answer
2021-11-22T11:31:38-0500

terms percentage = "\$3450 \times \frac{2}{10} \%"

terms percentage = "\$3450 \times \frac{2}{10 \times 100}=\$6.9"

payment=invoice - debt - terms percentage

=$3450 -$1500-$6.9

=$1943.1